OpenStudy (anonymous):

What's the answer to 7x=14

OpenStudy (anonymous):

2.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

or x=2

OpenStudy (anonymous):

Easy. Divide 7x from both sides. so 7x/7 = 14/7 would equal x=2

OpenStudy (anonymous):

7x=14 x=14/7 x=2